OpenStudy (anonymous):

plz help find the derivative of the fuction using thedefination of derivative state the domain of the function and the domain of its derivative f(x)=37 plz show me steps..

OpenStudy (anonymous):

f(x)=37? It's a constant..

OpenStudy (across):

\[ f'(x)=\lim_{\Delta x\to0}\frac{f(x+\Delta x)-f(x)}{\Delta x}. \]If \(f(x)=37\), then\[ f'(x)=\lim_{\Delta x\to0}\frac{37-37}{\Delta x}=0. \]The domain of both functions is \(\mathbb{R}\).